### v1

@@ 4 Error Decomposition
$\begin{cases}\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}},&\\ \lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}},&\end{cases}$ (4.4)
@@
×$\qquad\times\left(\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{4}\right)^{1/2}$
@@ 5.1 Bounding Approximation Error
$\leq\begin{cases}\kappa^{2(r+\alpha-1)},&\text{ when }r+\alpha\geq 1,\\ (r+\alpha)^{r+\alpha}(1-r-\alpha)^{1-r-\alpha}\lambda_{T}^{r+\alpha-1},&\text{when }r+\alpha<1,\end{cases}$
@@ 5.2 Bounding Initial Error
Let $10\leq\theta_{1}<$ 1, $10\leq\theta_{2}<$ 1, and $\eta_{t},\lambda_{t}$ be defined as (4.4). Then the following bounds hold for $1\leq l\leq m$ with $l\in\mathbb{N}$:
@@
=$\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(m+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(l+t_{0})^{1-\theta_{1}-\theta_{2}}\right],&\text{ when }\theta_{1}+\theta_{2}\not=1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{m+t_{0}+1}{t_{0}+l}\right),&\text{ when}\theta_{1}+\theta_{2}=1.\end{cases}$
@@
    $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, and $t_{0}>0$, corresponding to the online setting;
@@
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$. Then, for any $T\geq t_{0}+1$, $t_{0}\geq 0$, $10\leq\theta_{1}<$ 1, and $10\leq\theta_{2}<$ 1, the quantity
@@
$\leq c_{2}\bar{\eta}^{-2(r+\alpha)}\begin{cases}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})}\exp\{-\tau\bar{\eta}\bar{\lambda}(T+t_{0})^{1-\theta_{1}-\theta_{2}}\},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ (t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ (T+t_{0})^{-2(r+\alpha)(1-\theta_{1})},&\text{ when }\theta_{1}+\theta_{2}>1,\end{cases}$
@@ 5.3 Bounding Drift Error
In the finite-horizon setting, where $\lambda_{t}=\bar{\lambda}$ is fixed depending on $T$, we have $\mathcal{T}_{3}=0$. Therefore, it is sufficient to bound $\mathcal{T}_{3}$ under the regime of decaying step sizes and regularization parameters. In what follows, we focus on the setting where $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1.
@@
where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
@@
$\leq\|S^{\dagger}\|_{\mathrm{HS}}\bar{\lambda}\left\lvert(t+t_{0}-1)^{-\theta_{2}}-(t+t_{0})^{-\theta_{2}}\right\rvert\begin{cases}\kappa^{2r-2},&\text{ when }r\geq 1,\\ r^{r}(1-r)^{1-r}\lambda_{t}^{r-1},&\text{ when }r<1,\end{cases}$
@@
where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
@@
$\mathcal{T}_{3}=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
@@
$\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}\leq\begin{cases}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\},&\text{ when }\theta_{1}+\theta_{2}<1,\\ (t+t_{0})^{\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ 1,&\text{ when }\theta_{1}+\theta_{2}>1.\end{cases}$
@@
×$\times\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}.$
@@
$\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
@@
We will only use the bound for the case $\theta_{1}+\theta_{2}=1$ in the above proposition, as it provides better convergence rates than the other cases. Note that we cannot guarantee convergence of the estimation error when $\theta_{1}+\theta_{2}>1$. However, convergence of the prediction error is essured when $\theta_{1}+\theta_{2}>1$ with $10<\theta_{1},\theta_{2}<$ 1, and both the prediction and estimation errors converge when $\theta_{1}+\theta_{2}\leq 1$. The proofs for the remaining cases are similar and are omitted here to avoid repetition.
@@ 5.4 Bounding Sample Error
Suppose that Assumption 3 holds with $0<s\leq 1$. If $t_{0}\geq 1$, $T\geq t_{0}+1$, $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and the following condition holds for any $t\in\mathbb{N}_{T}$:
@@
×$\times\sum_{t=1}^{T}\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{\eta_{t}^{2}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{1+2\alpha-s}}.$
@@
The next proposition is used to bound $\mathcal{T}_{4}$ in the finite-horizon setting. We define $00^{0}:=$ 0 for convenience.
@@
Let $v\geq 0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, $10<\theta_{3}<$ 1, $\theta_{4}>0$, and $\eta_{1}$, $\lambda_{1}$ be constants independent of $T$. Then, there exists a constant $\delta_{3}$, independent of $T$, such that for any $T\geq 2$,
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\delta_{3}\begin{cases}T^{v\theta_{3}+(1-v)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }0\leq v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1.\end{cases}$
@@
We divide the proof into three cases: $v=0$, $v>0$, and $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$. The third case is an improvement upon the analysis in the second case.
@@
$\leq$ $1+\frac{1}{\bar{\eta}}+\frac{1}{\bar{\eta}}\begin{cases}\frac{\left(\bar{\eta}T\right)^{1-v}}{1-v},&\text{ when }0<v<1,\\ \log(\bar{\eta}T),&\text{ when }v=1,\\ \frac{1}{v-1},&\text{ when }v>1,\end{cases}$
$\leq$ $\delta_{3}\begin{cases}T^{1-v+\theta_{3}v},&\text{ when }0<v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1,\end{cases}$
@@
Case 3: $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$
@@
[…]
@@
[…]
@@
Suppose that Assumption 3 holds with $0<s\leq 1$. Let $\alpha\in[0,\frac{1}{2}]$, and set the parameters $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1 and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Additionally, assume that $T\geq 2$, $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and there exists a constant $\widetilde{M}$ independent of $T$, such that for all $t\in\mathbb{N}_{T}$,
@@
If $01+2\alpha-s>$ 0, applying Lemma 5.2 (3) for $11\leq t\leq T-$ 1, we obtain the following estimate, which also holds for $t=T$,
@@
If $01+2\alpha-s=$ 0, we derive the bound
@@ 5.5 Key Bounds for Estimating Prediction Error
$f(t+1):=$ $\mathcal{T}_{1}(t+1)+\mathcal{T}_{2}(t+1)+\mathcal{T}_{3}(t+1)$
@@ 6 Convergence Analysis in Expectation
$+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }s=1.\end{cases}$
@@
for any $T\geq t_{0}+1$, where the constant $c_{1,1}=c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is independent of $T$. Let $c_{1,1}$ be sufficiently large such that (6.1) holds true for $11\leq T<t_{0}+$ 1.
@@
for any $T\geq t_{0}+1$, where $c_{1,2}=c_{1}\bar{\lambda}^{2\min\{r,1\}}+c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is a constant independent of $T$. Let $c_{1,2}$ be sufficiently large such that (6.2) holds true for $11\leq T<t_{0}+$ 1.
@@
$+\widetilde{c_{4}}\begin{cases}T^{-\theta_{3}}\log T,&\text{ when}s=1,\\ T^{-\theta_{3}},&\text{ when }s<1.\end{cases}$
@@ 7 Convergence Analysis in High Probability
In this section, we derive the high-probability error bounds presented in Subsection 2.3. Our proofs are mainly based on the following proposition. This proposition is from [43, Proposition A.3], and is an extension of [39, Theorem 3.4].
@@
    ×$\times\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
@@
    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators and using Assumption 3 and (3) in Lemma 5.2, it follows that if $02\alpha+1-s>$ 0,
@@
    ×$\quad\quad\times\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\},$
    else if $02\alpha+1-s=$ 0,
@@
    ×$\times\frac{\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}.$
@@
    Case 1: If $12\alpha+1-s>$ 1, then
@@
    Case 2: If $12\alpha+1-s=$ 1, then
@@
    Case 3: If $10<2\alpha+1-s<$ 1, then
@@
    Case 4: If $02\alpha+1-s=$ 0. Since
@@
$\tau_{1}^{2}$ $=d_{5}\begin{cases}(T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s=1.\\ \end{cases}$
@@
×$\times\log^{2}\frac{2}{\delta}\begin{cases}1,&\text{ when }s<1,\\ \log(T+t_{0}),&\text{ when }s=1.\end{cases}$
@@
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,1}\begin{cases}(T+t_{0})^{-\theta_{1}}\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta},&s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{4}\frac{2}{\delta},&s=1.\\ \end{cases}$
@@
Since it is easy to verify that $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be bounded uniformly by a constant, we can choose $c_{2,1}$ to be sufficiently large such that the bound holds true for $11\leq T<t_{0}+$ 1.
@@
×$\left.\times\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$
@@
    If $12\min\{r,1\}+s<$ 1, i.e., $12r+s<$ 1, choose $\theta_{1}=\frac{1+2\min\{r,1\}+s}{3+2\min\{r,1\}+s}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{2}=\frac{2}{3+2\min\{r,1\}+s}=\frac{2}{3+2r+s}$. Then,
@@
Since $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be uniformly bounded by a constant, we can choose $c_{2,2}$ to be sufficiently large so that the bound also holds in this case.
@@
Suppose Assumption 3 and 5 hold. Let $\alpha\in[0,\frac{1}{2}]$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1, and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Assume $T\geq 2$ and $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$. Then,
@@
    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators, Assumption 3, and Lemma 5.2 (3), if $02\alpha+1-s>$ 0, then
@@
    ×$\times\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\},$
    and if $02\alpha+1-s=$ 0, then
@@
    ×$\times\begin{cases}T^{(1+2\alpha-s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
@@
    ×$\times\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s.\end{cases}$
@@
$\tau_{2}^{2}=d_{8}\delta_{3}\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$
@@
$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8(d_{7}+d_{8}\delta_{3})\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$ (7.32)
@@
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$

### v3

@@ 4 Error Decomposition
[…]
@@
$\qquad\times\left(\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{4}\right)^{1/2}$
@@ 5.1 Bounding Approximation Error
[…]
@@ 5.2 Bounding Initial Error
Let $0\leq\theta_{1}<1$, $0\leq\theta_{2}<1$, and $\eta_{t},\lambda_{t}$ be defined as (4.4). Then the following bounds hold for $1\leq l\leq m$ with $l\in\mathbb{N}$:
@@
$\sum_{t=l}^{m}\eta_{t}\lambda_{t}$ $=\bar{\eta}\bar{\lambda}\sum_{t=l}^{m}(t+t_{0})^{-\theta_{1}-\theta_{2}}\geq\bar{\eta}\bar{\lambda}\int_{l}^{m+1}(x+t_{0})^{-\theta_{1}-\theta_{2}}dx$
@@
    $0<\theta_{1}<1$, $0<\theta_{2}<1$, and $t_{0}>0$, corresponding to the online setting;
@@
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$. Then, for any $T\geq t_{0}+1$, $t_{0}\geq 0$, $0\leq\theta_{1}<1$, and $0\leq\theta_{2}<1$, the quantity
@@
[…]
@@ 5.3 Bounding Drift Error
In the finite-horizon setting, where $\lambda_{t}=\bar{\lambda}$ is fixed depending on $T$, we have $\mathcal{T}_{3}=0$. Therefore, it is sufficient to bound $\mathcal{T}_{3}$ under the regime of decaying step sizes and regularization parameters. In what follows, we focus on the setting where $0<\theta_{1}<1$ and $0<\theta_{2}<1$.
@@
where $\widetilde{c}_{3}$ is a constant independent of $_{0}, tt$, $\bar{\eta}$, and $\bar{\lambda}$.
@@
[…]
@@
where $\widetilde{c}_{3}$ is a constant independent of $_{0}, tt$, $\bar{\eta}$, and $\bar{\lambda}$.
@@
[…]
@@
[…]
@@
$\times\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}.$
@@
[…]
@@
We will only use the bound for the case $\theta_{1}+\theta_{2}=1$ in the above proposition, as it provides better convergence rates than the other cases. Note that we cannot guarantee convergence of the estimation error when $\theta_{1}+\theta_{2}>1$. However, convergence of the prediction error is essured when $\theta_{1}+\theta_{2}>1$ with $0<\theta_{1},\theta_{2}<1$, and both the prediction and estimation errors converge when $\theta_{1}+\theta_{2}\leq 1$. The proofs for the remaining cases are similar and are omitted here to avoid repetition.
@@ 5.4 Bounding Sample Error
Suppose that Assumption 3 holds with $0<s\leq 1$. If $t_{0}\geq 1$, $T\geq t_{0}+1$, $0<\theta_{1}<1$, $0<\theta_{2}<1$, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and the following condition holds for any $t\in\mathbb{N}_{T}$:
@@
$\times\sum_{t=1}^{T}\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{\eta_{t}^{2}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{1+2\alpha-s}}.$
@@
The next proposition is used to bound $\mathcal{T}_{4}$ in the finite-horizon setting. We define $0^{0}:=0$ for convenience.
@@
Let $v\geq 0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, $0<\theta_{3}<1$, $\theta_{4}>0$, and $\eta_{1}$, $\lambda_{1}$ be constants independent of $T$. Then, there exists a constant $\delta_{3}$, independent of $T$, such that for any $T\geq 2$,
[…]
@@
We divide the proof into three cases: $v=0$, $v>0$, and $0<v<1$ with $\theta_{3}+\theta_{4}<1$. The third case is an improvement upon the analysis in the second case.
@@
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq$ $1+\int_{0}^{T-1}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}dt\leq 1+\frac{1}{\bar{\eta}}\int_{0}^{\bar{\eta}(T-1)}\frac{1}{1+t^{v}}dt$ (5.14)
$\leq$ $1+\frac{1}{\bar{\eta}}\left(1+\int_{1}^{\bar{\eta}(T-1)}t^{-v}dt\right)$
@@
Case 3: $0<v<1$ with $\theta_{3}+\theta_{4}<1$
@@
$\mathcal{A}_{1}$ $\leq\int_{0}^{T^{\theta_{3}+\theta_{4}}}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}dt\leq\frac{1}{\bar{\eta}}+\frac{1}{\bar{\eta}}\int_{1}^{\bar{\eta}T^{\theta_{3}+\theta_{4}}}\frac{1}{t^{v}}dt$
@@
$\int_{k}^{+\infty}x^{-v}\exp\{-2\lambda_{1}\eta_{1}x\}dx\leq k^{-v}\frac{\exp\{-2\lambda_{1}\eta_{1}k\}}{2\lambda_{1}\eta_{1}}<\infty,$
@@
Suppose that Assumption 3 holds with $0<s\leq 1$. Let $\alpha\in[0,\frac{1}{2}]$, and set the parameters $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $0<\theta_{3}<1$ and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Additionally, assume that $T\geq 2$, $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and there exists a constant $\widetilde{M}$ independent of $T$, such that for all $t\in\mathbb{N}_{T}$,
@@
If $1+2\alpha-s>0$, applying Lemma 5.2 (3) for $1\leq t\leq T-1$, we obtain the following estimate, which also holds for $t=T$,
@@
If $1+2\alpha-s=0$, we derive the bound
@@ 5.5 Key Bounds for Estimating Prediction Error
$f(t+1)$= $\mathcal{T}_{1}(t+1)+\mathcal{T}_{2}(t+1)+\mathcal{T}_{3}(t+1)$
@@ 6 Convergence Analysis in Expectation
[…]
@@
for any $T\geq t_{0}+1$, where the constant $c_{1,1}=c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is independent of $T$. Let $c_{1,1}$ be sufficiently large such that (6.1) holds true for $1\leq T<t_{0}+1$.
@@
for any $T\geq t_{0}+1$, where $c_{1,2}=c_{1}\bar{\lambda}^{2\min\{r,1\}}+c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is a constant independent of $T$. Let $c_{1,2}$ be sufficiently large such that (6.2) holds true for $1\leq T<t_{0}+1$.
@@
[…]
@@ 7 Convergence Analysis in High Probability
In this section, we derive the high-probability error bounds presented in Subsection 2.3. Our proofs are mainly based on the following proposition. This proposition is from [40, Proposition A.3], and is an extension of [36, Theorem 3.4].
@@
    $\times\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
@@
    $\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators and using Assumption 3 and (3) in Lemma 5.2, it follows that if $2\alpha+1-s>0$,
@@
    $\quad\quad\times\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\},$
    else if $2\alpha+1-s=0$,
@@
    $\times\frac{\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}.$
@@
    Case 1: If $2\alpha+1-s>1$, then
@@
    Case 2: If $2\alpha+1-s=1$, then
@@
    Case 3: If $0<2\alpha+1-s<1$, then
@@
    Case 4: If $2\alpha+1-s=0$. Since
@@
[…]
@@
[…]
@@
[…]
@@
Since it is easy to verify that $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}$ for $1\leq T<t_{0}+1$ can be bounded uniformly by a constant, we can choose $c_{2,1}$ to be sufficiently large such that the bound holds true for $1\leq T<t_{0}+1$.
@@
$\left.\times\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$
@@
    If $2\min\{r,1\}+s<1$, i.e., $2r+s<1$, choose $\theta_{1}=\frac{1+2\min\{r,1\}+s}{3+2\min\{r,1\}+s}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{2}=\frac{2}{3+2\min\{r,1\}+s}=\frac{2}{3+2r+s}$. Then,
@@
Since $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ for $1\leq T<t_{0}+1$ can be uniformly bounded by a constant, we can choose $c_{2,2}$ to be sufficiently large so that the bound also holds in this case.
@@
Suppose Assumption 3 and 5 hold. Let $\alpha\in[0,\frac{1}{2}]$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $0<\theta_{3}<1$, and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Assume $T\geq 2$ and $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$. Then,
@@
    $\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators, Assumption 3, and Lemma 5.2 (3), if $2\alpha+1-s>0$, then
@@
    $\times\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\},$
    and if $2\alpha+1-s=0$, then
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]